OpenStudy (anonymous):

Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2.

OpenStudy (anonymous):

first find g(2)=-2(2)-3 that is -7 then find f(g(x))=f(g(2))=f(-7)=-(-7)+4 that is nothing but 11...

OpenStudy (anonymous):

ohh okay because i put the answer -7 but i totally forgot to solve f(g(x)) thank you so much :)

OpenStudy (anonymous):

@chetan552

OpenStudy (anonymous):

Mistakes makes us perfect....@anastasiad239